OpenStudy (anonymous):

I need help with this step by step Evaluate the indefinite integral. \int \frac{(\ln(x))^4}{x} dx by substitution

OpenStudy (anonymous):

\[\int\frac{\big(\ln x\big)^4}{x}~dx~~?\]

OpenStudy (anonymous):

If so, let \(u=\ln x\), then \(du=\dfrac{1}{x}~dx\): \[\int u^4~du\] Pretty simple from here.